OpenStudy (ajr262636):

7. Determine whether the graphs of the given equations are parallel, perpendicular, or neither Y = -2x + 3 2x + y = 7 A) Parallel B) Perpendicular C) Neither 8. Determine whether the graph of the given equations are parallel, perpendicular, or neither. Y = 4x – 2 -x + 4y = 0 A) Parallel B) perpendicular C) neither

OpenStudy (tkhunny):

What is your plan for doing that? If you put them in the Point-Slope form, it's pretty easy to tell. Actually Solving the system is another good way to go.

OpenStudy (ajr262636):

I Dont know how to do it.

OpenStudy (anonymous):

So these are your equations: y = -2x + 3 2x + y = 7 Now, when they are in the form y = mx + a, m:= the gradient, and figuring out the gradient is a good way to start this. So, converting your equations into the above format is a nice way to begin. The first one has already been done (y = -2x + 3), try and rearrange the other so that the y is on the one side

OpenStudy (tkhunny):

Absolutely unacceptable response. You MUST have SOMETHING. Plus, I just told you two ways to proceed. Give it a go before someone else does the work and you fail to learn anything.

OpenStudy (ajr262636):

heres my problem! This is my second time taking algebra 1 and when i look at it it just looks like a bunch of letters and numbers jumbled together and it makes no sence, i am trying to learn but i havent been able to. i dont want the answers, i want someone to dumb it down for me so i can understand it and do it myself

OpenStudy (tkhunny):

Why do you think it can be "dumbed down"? It's supposed to take thought. That's part of the purpose of the study of mathematics. It would help immensely if you would answer questions directly, rather than making excuses why you have no questions. Have you ever heard of the Point-Slope form? Can you solve an equation for a given variable? Have you ever seen a solution to a linear system? If you've already been all the way through Algebra I, what do you remember? There has to be something. You will learn by addressing it, not by making excuses for it.

OpenStudy (ajr262636):

yes i have heard of a point slope, i dont know how to graph it but i am able to find it when it has already been graphed. no i dont know how to solve for an equation for a given varaible, i dont even know what that is. what is a solution for a linear system? I dont remember much because i live in mexico and i went to a public school so all of my classes were in spanish... i do speak spanish but i learned it is spanish so its hard for me to be able to translate everything i learned in english, and what i learned and remember i have no troubles with.

OpenStudy (tkhunny):



OpenStudy (anonymous):

Don't worry about it - you probably know more than you think. A lot of places use different words or take different ways of approaching problems, you just need to understand it for yourself. Take it slowly, step by step. I'm from the UK, and I've never heard it referred to as a Point-Slope but that doesn't mean I can't do it.

OpenStudy (tkhunny):

Perfect. You need to find the slope. Most commonly, you solve for y. y = -2x + 3 This is already in Slope-Intercept form. Just read off the slope. Slope = -2 2x + y = 7 This is not quite in Slope-Intercept form. Can you get it there?

OpenStudy (ajr262636):

thanks, i get the parallel thing but i dont know how to graph it or know with just the equation whether it is parallel or perpendicular

OpenStudy (tkhunny):

Okay, now answer my question.

OpenStudy (tkhunny):

@livias.random What is it in the UK from your studies? Slope-Intercept: y = mx + b Point-Slope: (y-y0)=m(x-x0)

OpenStudy (ajr262636):

*scrathes head* do i have to solve the equation????? what i want keep doing is solve for the unknown, but i dont think thats right

OpenStudy (tkhunny):

Solve for y. This will give y = mx + b. That 'm' is the slope. Works every time that it works.

OpenStudy (anonymous):

@tkhunny Yeah so we don't make such a big deal of the name but probably just the standard form of the linear equation? I just call it mx + c because I'm lazy. And with 'Point-Slope: (y-y0)=m(x-x0)' Instead of y0 we would use b. (That could just be an SQA thing though, don't know the English curriculum very well)

OpenStudy (ajr262636):

im lost.........

OpenStudy (tkhunny):

Fair enough. Thanks. @ajr262636 2x + y = 7 You MUST have the skill to solve this for anything required. Solve for y by subtracting 2x. Please demonstrate.

OpenStudy (ajr262636):

oh subtracting 1 sec

OpenStudy (tkhunny):

That is division by 2 - sort of. That is not subtracting 2x. Please subtract 2x, instead.

OpenStudy (ajr262636):